# On Prompt-Driven Safeguarding for Large Language Models

### v2

@@ 4.3 Robustness Analysis
In DRO’s anchoring process (§3.1), we use a set of anchor data to derive the low-dimensional representation space and refusal direction. We are interested in how robust DRO is to the choices of anchor data. We conduct ablation study for anchor data from the two aspects that compose the anchor data. For queries that were originally collected with careful controls (§2.1; used in §2.3 and main experiments), we keep the 100 synthetic harmless ones but replace the 100 synthetic harmful ones with the 100 queries from AdvBench. Note that these queries (after replacement) are also used for the subsequent DRO training. This replacement leads to the format gap between the harmless and the new harmful queries, i.e., the former are all “How to do” questions while the latter are all imperative “Do something” instructions, which simulates the case where the queries are collected with less careful controls. For basic safety prompts, we originally equipped queries with all three basic safety prompts (default, mistral, and short; $k=3$) to form eight groups of data points for anchoring ($2\times(1+k)$; §3.1), and then optimized the three different basic safety prompts separately (§4.1). Now we use only the default one ($k=1$) to form four groups of data points for anchoring, but then optimize the short one, which results in a gap between the basic safety prompt used for anchoring (default) and the one to be optimized (short). This enables us to fairly assess whether using a single safety prompt can still effectively anchor a low-dimensional space that captures the features related to models’ refusal behaviors.
@@
The results of ablation study for anchor data are shown in Table 4. We find that DRO still notably enhances the safeguarding performance. However, when the queries are less carefully controlled, the general model capability can be slightly degraded (59.0 vs. 63.5). It is probably due to the distraction of the spurious features that can be used to distinguish harmful and harmless queries, such as the textual format. We also observe that when we use only a single safety prompt for anchoring, the safeguarding performance is slightly inferior to that when we use multiple ones (4.1 vs. 2.3). It suggests that a single safety prompt may introduce biases that hinder accurately capturing the most salient features related to models’ refusal behaviors. But overall, DRO exhibits reasonable robustness to the choices of anchor data, and we suggest applying proper query controls and combining multiple basic safety prompts for the anchor data to achieve better safeguarding performance.
4.4 Interpretability Analysis
5 Related Work
Large Language Model Safety
Research on LLM safety aims to avoid LLMs producing contents that may cause harm to individuals and society. Previous work extensively studied to eliminate undesirable attributes from LLM-generated texts, such as toxic language and hate speech (Xu et al., 2020; Sun et al., 2022; Adolphs et al., 2023; Zheng et al., 2023). As the capabilities of LLMs keep growing, researchers are paying increasing attention to preventing LLMs from assisting queries or instructions with harmful intents, i.e., training or teaching them to refuse (Shaikh et al., 2023; Touvron et al., 2023; Bai et al., 2022a, b; OpenAI, 2023), which is the focus of our work. Recent work has also noticed the more complex jailbreak attacks, which manipulate LLMs into providing assistance by obfuscating LLMs’ recognition of the queries’ harmfulness (Zou et al., 2023; Wei et al., 2023; Zeng et al., 2024). Our work may inspire future research to delve into the intrinsic causes of LLMs’ vulnerabilities and stimulate more comprehensive and principled safeguarding methods.
@@ 6 Conclusion
We investigate the working mechanisms of safety prompts in safeguarding LLMs from the perspective of model representations. We find that safety prompts may not improve LLMs in recognizing the harmfulness of queries, but rather increase LLMs’ overall probability of refusing queries by moving queries’ representations in a “higher-refusal” direction. Inspired by this, our proposed DRO method optimizes continuous safety prompts by moving queries’ representations in the low-dimensional space along or opposite the estimated refusal direction, in which the model’s refusal probability increases. We show that DRO brings remarkable improvement in safeguarding performance, does not compromise the general model capability, and exhibits reasonable robustness to the choices of the data used for anchoring the low-dimensional space. We hope the empirical analysis and the proposed methodology in this work can inspire future research on LLM safety.
Impact Statements
@@
We have released the data, code, and experimental results at https://github.com/chujiezheng/LLM-Safeguard to facilitate reproducible research. Our experiments were run on A100 40GB GPUs. The open-source models and data used in this work are listed in Appendix A.
@@ Appendix B Basic Safety Prompts Used in Experiments
default-only normal-$\to$ short
@@
default-only normal-$\to$ short
@@
default
You are a helpful, respectfuliNdEx honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content. Please ensure that your responses are socially unbiased and positive in nature.
@@
default
You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should have include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content febbra Please ensure that your responses are socially unbiased and positive in nature.
@@
default
You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content. Please ensure that your responses[s] socially unbiased and positive in nature.
@@
mistral
You are an AI assistant. Always assist with care, respect, and truth. Please respond with utmost utility yet securely and avoid harmful, unethical, prejudiced, or negative content. Also ensure replies promote[s]ness and positivity.
@@
short
You are a helpful, respectful and honest assistant. Always answer as helpfully[s] possible, while being safe.

### v3

@@ 4.4 Robustness Analysis
In DRO’s anchoring process (§ 3.1), we use a set of anchor data to derive the low-dimensional representation space and refusal direction. We are interested in how robust DRO is to the choices of anchor data. We conduct ablation study for anchor data from the two aspects that compose the anchor data. For queries that were originally collected with careful controls (§ 2.1; used in § 2.3 and § 4.1), we keep the 100 synthetic harmless ones but replace the 100 synthetic harmful ones with the 100 queries from AdvBench. Note that these queries (after replacement) are also used for the subsequent DRO training. This replacement leads to the format gap between the harmless and the new harmful queries, i.e., the former are all “How to do” questions while the latter are all “Do something” instructions, which simulates the case where the queries are collected with less careful controls. For basic safety prompts, we originally equipped queries with all three basic safety prompts (default, mistral, and short; $k=3$) to form eight groups of data points for anchoring ($2\times(1+k)$; § 3.1), and then optimized the three different basic safety prompts separately (§ 4.1). Now we use only the default one ($k=1$) to form four groups of data points for anchoring, but then optimize the short one, which results in a gap between the basic safety prompt used for anchoring (default) and the one to be optimized (short). This enables us to fairly assess whether using a single safety prompt can still anchor a low-dimensional space that captures the features related to models’ refusal behaviors.
@@
The results of ablation study for anchor data are shown in Table 4. We find that DRO still notably enhances the safeguarding performance. However, when the queries are less carefully controlled, the models’ general performance can be slightly degraded (59.0 vs. 63.5). It is probably due to the distraction of the spurious features that can be used to distinguish harmful and harmless queries, such as the textual format. We also observe that when we use only a single safety prompt for anchoring, the safeguarding performance is slightly inferior to that when we use multiple ones (4.1 vs. 2.3). It suggests that a single safety prompt may introduce biases that hinder accurately capturing the most salient features related to models’ refusal behaviors. But overall, DRO exhibits reasonable robustness to the choices of anchor data, and we suggest applying proper query controls and combining multiple basic safety prompts for the anchor data to achieve better safeguarding performance.
@@
4.5 Interpretability Analysis
5 Related Work
Large Language Model Safety
Research on LLM safety aims to avoid LLMs producing contents that may cause harm to individuals and society. Previous work extensively studied to eliminate undesirable attributes from LLM-generated texts, such as biases, toxic language, and hate speech (Xu et al., 2020; Sun et al., 2022; Adolphs et al., 2023; Zheng et al., 2023, 2024b). As the capabilities of LLMs keep growing, researchers are paying increasing attention to preventing LLMs from assisting queries or instructions with harmful intents, i.e., training or teaching them to refuse (Shaikh et al., 2023; Touvron et al., 2023; Bai et al., 2022a, b; OpenAI, 2023), which is the focus of our work. Recent work has also noticed the more complex jailbreak attacks, which manipulate LLMs into providing assistance by obfuscating LLMs’ recognition of the queries’ harmfulness (Zou et al., 2023; Wei et al., 2023; Liu et al., 2024; Zeng et al., 2024). Our work can inspire future research to delve into the intrinsic causes of LLMs’ vulnerabilities and stimulate more principled safeguarding methods.
@@ 6 Conclusion
We investigate the working mechanisms of safety prompts in safeguarding LLMs from the perspective of model representations. We find that safety prompts do not clearly improve LLMs in recognizing the harmfulness of queries, but rather increase LLMs’ overall probability of refusing queries by moving queries’ representations in a “higher-refusal” direction. Drawing this inspiration, our proposed DRO method optimizes continuous safety prompts by moving queries’ representations in the low-dimensional space along or opposite the estimated refusal direction, in which the model’s refusal probability increases. We show that DRO brings remarkable improvement in safeguarding performance on both out-of-domain and jailbreak benchmarks, does not compromise the models’ general performance, and exhibits reasonable robustness to the choices of the data used for anchoring the low-dimensional space. We hope the empirical analysis and the proposed methodology in this work can inspire future research on LLM safety.
Impact Statement
@@
Acknowledgments
We thank Yufei Tian, Rohan Wadhawan, Yu (Bryan) Zhou, Haw-Shiuan Chang, Po-Nien Kung, and other members of the UCLA PlusLab & NLP group for their constructive feedback and discussions. We thank Xiaogeng Liu and Nan Xu for sharing the GCG jailbreak prompts.
This work was supported by the National Science Foundation for Distinguished Young Scholars (with No. 62125604). This work was also supported by the NSFC projects (Key project with No. 61936010).
@@ Appendix B Basic Safety Prompts Used in Experiments
default-only $\to$ short
@@
default-only $\to$ short
@@
default You are a helpful, respectfuliNdEx honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content. Please ensure that your responses are socially unbiased and positive in nature.
@@
default You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should have include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content febbra Please ensure that your responses are socially unbiased and positive in nature.
@@
default You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content. Please ensure that your responses[s] socially unbiased and positive in nature.
@@
mistral You are an AI assistant. Always assist with care, respect, and truth. Please respond with utmost utility yet securely and avoid harmful, unethical, prejudiced, or negative content. Also ensure replies promote[s]ness and positivity.
@@
short You are a helpful, respectful and honest assistant. Always answer as helpfully[s] possible, while being safe.